OpenStudy (anonymous):

tanx/2=tanx/secx+1 Need help.

OpenStudy (aum):

\[ \cos(2x) = 2\cos^2(x) - 1 \\ \cos^2(x) = \frac 12(1 + \cos(2x)) \\ \text{ } \\ \cos(2x) = 1 - 2\sin^2(x) \\ \sin^2(x) = \frac 12(1 - \cos(2x)) \\ \tan^2(x) = \frac{\sin^2(x)}{\cos^2(x)} = \frac{1-\cos(2x)}{1+\cos(2x)} \\ \tan^2\left(\frac x2\right) = \frac{1-\cos(x)}{1+\cos(x)} = \frac{1-\cos(x)}{1+\cos(x)}*\frac{1+\cos(x)}{1+\cos(x)} = \frac{1-\cos^2(x)}{(1+\cos(x))^2} \\ \text{ } \\ \tan^2\left(\frac x2\right) = \frac{\sin^2(x)}{(1+\cos(x))^2} \\ \tan\left(\frac x2\right) = \frac{\sin(x)}{1+\cos(x)} = \large \frac{\frac{\sin(x)}{\cos(x)}}{\frac{1+\cos(x)}{\cos(x)}} = \normalsize \frac{\tan(x)}{\sec(x)+1} \]